OpenStudy (anonymous):

Help please; Integrate following : 1/(e^x + 1) Thanks in advance.

OpenStudy (anonymous):

Rewrite it as: \[\int\limits \frac{1+e^x-e^x}{e^x+1} dx=\int\limits \frac{e^x+1}{e^x+1}dx-\int\limits \frac{e^x}{e^x+1} dx\]. The first integral reduces to the integral of 1. The second is a u-sub.

OpenStudy (anonymous):

ln(e^x(e^x+1))+c

OpenStudy (anonymous):

Rewrite it as: \[\int\limits \frac{1+e^x-e^x}{e^x+1} dx=\int\limits \frac{e^x+1}{e^x+1}dx-\int\limits \frac{e^x}{e^x+1} dx\]. The first integral reduces to the integral of 1. The second is a u-sub.

OpenStudy (anonymous):

let p = e^x+1. So integrand becomes 1/p. Integral is ln(p) + c. is this right?

OpenStudy (anonymous):

no. it seems wrong.

OpenStudy (anonymous):

Let e^x=t e^x dx = dt dx=dt/e^x=dt/t Hence the integral is dt/(t(t+1)) =dt/(t^2+t)=dt/((t+1/2)^2-1/4) let t+1/2=k the integral now is dk/(k^2-1/4) =ln(k-1/2)-ln(k+1/2)=ln(t)-ln(t+1)=ln(e^x)-ln(e^x+1)= x-ln(e^x+1)

OpenStudy (anonymous):

Damn shankvee.

OpenStudy (anonymous):

pls wait. let me go thru it.

OpenStudy (anonymous):

kiranbandal whenever you substitute you have to care care of dx. in your case dp=e^x dx

OpenStudy (anonymous):

malevolence what did i do? I just gave a method to integrate the function which is a pretty simple one.

OpenStudy (anonymous):

I'm just saying :P double substitution? haha.

OpenStudy (anonymous):

Shankvee you changed the variables and did more than one substitution ... That was way more complicated than it needed to be

OpenStudy (anonymous):

and Kiranbandal, you can also multiply top and bottom by e^(-x) it will also simplify it into less work than my previous post.

OpenStudy (anonymous):

You still get the answer dont you? There is not much complixation You should get used to it if you practice.

OpenStudy (anonymous):

ok, let me try "multiply top and bottom by e^(-x)"

OpenStudy (anonymous):

I mean, I'm not knocking your approach. Just seems like a lot work xP

OpenStudy (anonymous):

malevolence what is your problem with double substitutions? :D

OpenStudy (anonymous):

(e^x+1) e^-x ... what would it be?

OpenStudy (anonymous):

I suggest complicating it more by integrating 1/(t(t+1)) by parital fractions :P

OpenStudy (anonymous):

ok. i got it. thanks shankvee and malevolence19.

OpenStudy (anonymous):

my answer is: - ln | e^-x + 1 | + C. I divided top and bottom by e^x. So, i got f'(x) / f(x).

OpenStudy (anonymous):

actually when you nultiply and divide e^x you will have e^x in the denominator also so differential of e^x(1+e^x) is not equal to e^x

OpenStudy (anonymous):

so, is my answer wrong?

OpenStudy (anonymous):

unfortunately yes. :D

OpenStudy (anonymous):

see, d/dx of (e^-x + 1) is -e^x. so, i introduced -1 inside and outside of integrand.

OpenStudy (anonymous):

Shank-I don't have a problem with it :P Just not good for rookies xDDD